OpenStudy (anonymous):

given sin A =-4/5 and pi < A < 3pi/2 and sin B = 12/13 and pi/2 < B < pi find sin(a-b) help plz

OpenStudy (anonymous):

\[\sin(A-B)=\sin(A)\cos(B)-\cos(A)\sin(B)\]

OpenStudy (anonymous):

and you have sin(A) and sin(B) you need to find cos(A) and cos(B)

OpenStudy (anonymous):

\[\sin ^{2}A+\cos ^{2}A=1\] \[\sin ^{2}B+\cos ^{2}B=1\]